OpenStudy (anonymous):

Solve the equation. x/9=12+x/3

OpenStudy (anonymous):

the LCD=9 x/9=(108+3x)/9 x=108+3x -2x=108 x=-54

OpenStudy (anonymous):

thank you thank you thank you!!!

OpenStudy (anonymous):

You are welcome